Circular No. 6187

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1995 O1 Independent reports of the visual discoveries of a new comet have been

received from Alan Hale and Thomas Bopp, with available observations given below. All observers note the comet to be diffuse with some condensation and no tail, motion toward the west-northwest.

1995 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 23.264 18^{h}44.^{m}23 −32^{o}13.^{′}6 10.5 Hale
23.30 18 46 −32.6 10.8 Stevens
23.375 18 44.17 −32 11.^{′}2  Hale

A. Hale (Cloudcroft, NM). 0.41-m reflector. J. Stevens and T. Bopp (near Stanfield, AZ). 0.44-m f/4.5 Dobsonian re-

flector. Comet found while observing M70.

PSR J0538+2817 X. Sun, Max-Planck-Institut für Extraterrestrische Physik (MPIEP)

and Institute of High Energy Physics, Beijing; B. Aschenbach, W. Becker, and J. Truemper, MPIEP; and S. Anderson and A. Wolszczan, Pennsylvania State University, report the discovery of x-ray emission from the radio pulsar PSR J0538+2817 in S147 (= G180.0–1.7; cf. IAUC 6012) at count rates of 0.048 ± 0.013 and 0.051 ± 0.012 count/s in the broad (0.8–2.0 keV) and hard (0.5–2.0 keV) energy bands, respectively: “The x-ray source was detected during the ROSAT All Sky Survey at α = 5^{h}38^{m}26^{s}.0, δ = +28^{o}17′15′′ (equinox 2000.0), which is only 14′′ away from the radio pulsar (Anderson et al. 1995, work in progress), well within the error of the ROSAT PSPC position. There is no other object in the vicinity of the source in the SIMBAD and GSC databases. The short survey exposure time does not permit a timing analysis or a search for a compact nebula emission, although a large complex diffuse emission associated with SNR S147 was detected in the hard-energy band. Assuming a blackbody spectrum and a 10-km-radius emission area, and adopting the column density (1.2 × 10^{21}) and distance (1.5 kpc) derived from the radio-dispersion measure, the energy distribution of photons can be used to constrain the temperature and luminosity to ∼ 0.1 keV and 3 × 10^{32} erg/s. This temperature is in rough agreement with the prediction of standard neutron star cooling models using the spin-down age of 6 × 10^{5} yr.”

1995 July 23 Daniel W. E. Green